# ForgeHLS: A Large-Scale, Open-Source Dataset for High-Level Synthesis

### v2

@@ III-A Dataset Overview
Specifically, the data construction flow involves two key stages: collection and generation of kernels (Sec. III-B) and design space exploration for pragma insertion (Sec. III-C). Based on this flow, we can construct a dataset that is both comprehensive and versatile, suitable for multiple HLS-related downstream tasks (Sec. IV).
III-B Collection and Generation of Kernels
We adopt a dual approach: aggregating real-world HLS kernels from open-source repositories and generating diverse algorithmic implementations through GPT-4o to collect rich kernels. For real-world HLS code, we aggregate data from a variety of established benchmarks to ensure broad coverage and applicability. These sources include: PolyBench [12], MachSuite [14], CHStone [13], and Rosetta [15]. Additionally, we include a significant collection from Vitis HLS official codebase [11].
For generated C++ code by GPT-4o, we provide a cross-verification workflow to ensure correctness, as shown in Fig. 2. We prioritize the inclusion of a wide range of HLS-compatible algorithms. ForgeHLS covers various algorithms: 1) fundamental operators, such as adders, counters, and dividers, which are critical for low-level hardware designs; 2) typical HLS computation-intensive algorithms, such as matrix multiplication, FFT, signal processing algorithms, I2C master/slave controllers, and IP checksum calculation, that are feasible for hardware implementation and commonly optimized in HLS. This category also includes commonly used kernels from deep learning, such as LSTM cells and DNN layers; 3) Leetcode algorithms, including representative C/C++ algorithms such as binary search and merge sort, solving classical algorithmic problems that are often used for algorithmic training and benchmarking; 4) RTL (Register-Transfer Level) designs implemented in HLS C code, including RTL chips, RTL IPs and RTL modules, following the same description and specification as those of the RTL dataset [25].
We input the specifications for the algorithm description and HLS code rules into GPT-4o, which then automatically generate the corresponding code without HLS pragmas. The HLS code rules ensure correct input and output parameter usage, maintain scalability for future pragma additions, enforce strict compliance with the C++11 standard, and prohibit the use of unsupported C/C++ constructs. The unsupported C/C++ constructs include system calls, dynamic memory usage, pointer limitations, recursive functions, undefined behaviors, and virtual functions and pointers. We run a single-pass HLS flow to validate the syntactic accuracy of LLM generated code. To validate the functional accuracy, we employ a cross-verification approach leveraging GPT-4o as an assistant. Specifically, we provided GPT-4o with an algorithm description and the corresponding implementation, requesting it to generate a Python version of the same algorithm. This Python implementation served as a reference model for validation. Only when the synthetic code meets both syntactic and functional correctness criteria, we collect it in our ForgeHLS. After that, we apply DSE on kernels to further enhance our dataset.
III-C Design Space Exploration for Pragma Insertion
Following this, we construct a pragma insertion tree to guide the selection of pragma factors. Specifically, we determine whether to enable or disable function inlining and pipelining on the first line inside each function. For the array partition pragma, we choose the target array, the dimension to apply, the partition factor, and the partition type. For loop optimizations, we decide whether to apply loop pipeline, whether to apply unroll, and if unrolling is applied, what unroll factor to choose.
It is worth emphasizing that the full DSE does not exhaustively enumerate all pragma permutations. Instead, it follows a set of essential HLS design principles: 1) if the outer loop of a nested loop structure is pipelined, the inner loop will not use the unroll pragma; 2) unroll and partition factor selection is restricted to powers of 2 within the range of the upper limit of the loop trip count or array size; 3) if the current branch of the pragma design tree becomes excessively large, pruning is performed to retain only the designs whose unroll and partition factors are equivalent. By applying these rules, we eliminate invalid pragma combinations, which helps in reducing the dataset size without compromising its quality. A traversal in design tree from the root node to a leaf node represents one pragma configuration. Each path encodes a sequence of pragma decisions, and the resulting leaf node corresponds to a unique design instance. This tree-based representation enables systematic exploration of the design space, where constraints can be enforced dynamically during traversal to avoid infeasible or redundant configurations.
However, due to the combinatorial nature of pragma factor selection, the number of design instances in full DSE grows exponentially with the number of applied pragmas. For complex kernels with more than 10 pragmas, full design space traversal can easily yield millions of design candidates. Performing HLS synthesis on such a large number of configurations is computationally prohibitive and contradicts the core objective of our dataset—efficiently generating high-quality, diverse design instances.
To address this challenge, we propose a Bayesian DSE strategy tailored for complex design cases. Instead of exhaustively evaluating the entire design space, Bayesian DSE leverages a probabilistic surrogate model to iteratively select promising configurations based on previous evaluation results. Algorithm 1 presents our design space exploration strategy based on Bayesian optimization. As a lightweight and flexible exploration technique, it aims to efficiently discover diverse high-quality design candidates without exhaustively enumerating the entire pragma space. Given a C/C++ kernel , we first extract its pragma configuration space using Clang static analysis (line 1). The algorithm begins by seeding the design set with a set of randomly sampled initial designs (lines 3 to 5), which serve as starting points for exploration. Then, for each optimization round, the algorithm iteratively performs Bayesian optimization over steps (lines 6 to 11). At each iteration, it generates a new design with pragma from the current pragma configuration (line 7), evaluates the design using HLS synthesis (line 8), and collects resource and latency information. Each is stored in (line 9). Based on the observed quality metrics, a cost is calculated (line 10) to reflect the overall design objective. Finally, the Bayesian optimizer updates its surrogate model and acquisition function based on the new observation, and selects the next pragma configuration accordingly. The algorithm continues until the optimization round completes.
This iterative process allows the exploration to focus on promising regions of the design space while maintaining diversity through random initialization. The accumulated design set , containing pragmas and QoR results, is returned for constructing the ForgeHLS dataset. Notably, our goal with DSE is to expand design space coverage rather than solely focus on identifying a Pareto curve. For downstream tasks such as QoR prediction, the value of non-Pareto-optimal designs in ForgeHLS lies in the trade-offs they represent, which provide valuable insights. Our code analysis extracts all relevant details about arrays and loops, and we randomly configure all the pragmas, running HLS to sample non-Pareto-optimal designs. These random starting points also serve as a source of data. After collecting all designs for a given kernel, we label the generated pareto points as Pareto within our dataset. Although these points may slightly deviate from ground-truth pareto-optimal designs as exhaustive evaluation of the entire design space is computationally infeasible, it is acceptable to use these near-optimal designs for model training and development. Users can also adopt their customized DSE algorithms to generate other near-optimal points. A more detailed analysis of the design space coverage in ForgeHLS will be provided in Sec. IV-A.
IV Evaluation
We utilize Vitis HLS 2023.2 to collect ground-truth latency and resource usage as our dataset labels. The clock is set to 10 ns and the target FPGA is xcu280-fsvh2892-2L-e.
IV-A Design Space Coverage
In this section, we evaluate the effectiveness of our approach in expanding the design space coverage, particularly with respect to latency and resource usage, which are essential metrics for model training in downstream tasks such as QoR prediction. For latency, we use the Worst Case Latency extracted from the Vitis HLS report, and for resource usage, we compute the Average Resource Usage (ARU) across four key resource (BRAM, FF, LUT, and DSP). The ARU is calculated using the following equation 1. represents the number of resources for which .
Progressive Dataset Expansion. We have progressively expanded our dataset to achieve broader kernel coverage. Initially, we used standard HLS benchmark suites, such as CHStone, MachSuite, PolyBench, and Rosetta. We then extended the dataset by incorporating kernels from the Vitis Examples industry database, which significantly improved coverage. Furthermore, the inclusion of kernels generated by LLM Synthetic Code contributed to another substantial increase in design space coverage. As shown in Fig. 5, these kernels, sourced from diverse origins, complement each other in expanding the design space and enriching dataset diversity.
Detailed Distribution of ForgeHLS. The detailed distribution of the dataset across different sources is shown in Fig. 4. We present key metrics of our dataset, including the code length, latency, number of pragmas and resource usage. Each data point in these plots represents an individual design, demonstrating the spread of design characteristics across sources. This analysis shows the incremental improvement in coverage as new sources are added, particularly highlighting how LLM-generated synthetic code helps fill gaps in kernel diversity and design complexity. From this distribution, it can be observed that the synthetic code exhibits a relatively rich distribution overall. While the number of pragmas and tokens in synthetic code is slightly lower compared to real-world benchmarks, the feature distribution in terms of resource usage and latency appears comparable. This indicates that its diversity in resource and latency characteristics is sufficient for robust model training.
Comparison with Existing Datasets. While we compare our dataset to HLSyn and DB4HLS in Fig. 5, it is important to note that we have not included comparisons with HLSDataset and HLSFactory. This omission is due to the incomplete accessibility of latency and resource data from these sources. Both HLSDataset and HLSFactory have only released a limited portion of their latency and resource metrics.
It is also worth noting that HLSyn’s design leverages AutoDSE [5], resulting in designs that tend to exhibit low resource usage. This leads to a lack of diversity in design solutions. For robust model training, it is crucial to consider not only Pareto-optimal solutions but also non-Pareto solutions. These non-Pareto solutions can also provide valuable insights into the trade-offs between resource usage and latency, contributing to the robustness of the model and its ability to generalize to a broader range of design scenarios.
Despite this limitation, as shown in Fig. 5, our dataset demonstrates superior coverage in terms of both latency and resource usage compared to HLSyn and DB4HLS. Our dataset’s larger scale and more comprehensive distribution make it better suited for addressing the diverse challenges encountered in real-world applications. This broader coverage enhances the robustness of model training for tasks like QoR prediction and hardware design optimization.
IV-B Downstream Task 1: HLS QoR Prediction
| MAPE (%) | Standard Test Set | Unseen Test Set | ||||
| Model-Trainset | DSP | FF | LUT | DSP | FF | LUT |
| RGCN-#[2] | 18.8 | 26.5 | 32.4 | 2191.7 | 1004.0 | |
| RGCN-#P-ForgeHLS | 47.4 | 25.4 | 21.8 | 251.7 | 1521.4 | 174.1 |
| RGCN-ForgeHLS | 84.7 | 228.4 | 44.1 | 123.1 | 511.7 | 62.5 |
| SAGE-#[2] | 17.0 | 20.8 | 18.7 | 136.2 | 818.2 | 103.5 |
| SAGE-#P-ForgeHLS | 52.8 | 26.1 | 12.7 | 219.0 | 2026.2 | 384.9 |
| SAGE-ForgeHLS | 87.1 | 101.6 | 34.3 | 94.1 | 1833.4 | 68.8 |
| LLaMA3-7B-Origin | 140.9 | 2448.5 | 217.1 | 142.9 | 2428.6 | 221.3 |
| LLaMA3-7B-DB4HLS | 72.4 | 901.7 | 602.4 | 70.2 | 3529.3 | 1529.5 |
| LLaMA3-7B-ForgeHLS | 23.1 | 61.7 | 26.0 | 55.1 | 111.2 | 59.7 |
| Qwen-7B-Origin | 1990.8 | 2250.6 | 309.8 | 2644.4 | 3791.5 | 413.7 |
| Qwen-7B-DB4HLS | 81.1 | 481.0 | 394.8 | 73.2 | 2665.7 | 951.6 |
| Qwen-7B-ForgeHLS | 46.6 | 49.3 | 27.0 | 84.3 | 96.1 | 65.0 |
| Mis-7B-Origin | 287.5 | 338.4 | 209.8 | 70.6 | ||
| Mis-7B-DB4HLS | 3369.1 | 2933.7 | 3780.8 | 3656.4 | ||
| Mis-7B-ForgeHLS | 51.8 | 54.5 | 22.9 | 58.6 | 73.3 | 37.1 |
@@ IV-B Downstream Task 1: HLS QoR Prediction
In this downstream task, we train GNN models and LLMs on various datasets for post-HLS QoR prediction. Specifically, we aim to explore two key questions: 1) can GNNs accurately predict the true QoR values in a dataset as large as ours? 2) can LLMs predict the corresponding QoR metrics after fine-tuning, and how accurate are their predictions?
To fairly compare the model’s generalization ability, we propose the Unseen Test Set, which is constructed by selecting 10% of kernels from ForgeHLS, resulting in a set of 57 kernels, each with its corresponding designs. By doing so, we ensure that all training sets do not include any kernels from the Unseen Test Set, allowing us to evaluate the model with on a consistent metrics. We also build Standard Test Set for each dataset. The Standard Test Set is created by performing a 90:10 split from each dataset designs, representing the model’s fitting ability. For the GNN training baseline from [2], this involves replicating the original methodology by training on the provided training set and evaluating the model using the corresponding test set. For GNN training on ForgeHLS and all LLMs training, the Standard Test Set is created by splitting the designs in a 90:10 ratio from the respective training sets. Unlike the Unseen Test Set, the Standard Test Set may contain kernels that overlap with those seen during training. It is important to note that the training set used for the Standard Test Set in ForgeHLS is the remaining kernel portion of the dataset not included in the Unseen Test Set. Therefore, the Unseen Test Set does not have any bias towards ForgeHLS.
We follow the knowledge rich approach from [2] for GNN training, where ADB files are compiled into control data flow graphs to serve as input. Each GNN is set as 5 layers with a hidden-dimension size of 300, and 300 epochs for Adam optimizer. For LLMs, we train mainstream open-source models by LoRA [27] fine-tuning the LLMs, with Code-QoR pairs as input in the format {"input": "code with HLS pragma", "output": {"lut": lut_num, "dsp": dsp_num, .%..}}. This setup enables the LLMs to directly predict QoR metrics. The key LoRA training settings are as follows: the LoRA rank is set to 32, and the LoRA alpha is set to 64. We use MAPE (Mean Absolute Percentage Error) as the metric to evaluate prediction errors. In the evaluation of QoR, we focus on LUT, DSP, and FF as the primary resource metrics. BRAM resources were not included in the evaluation because our experiments primarily target designs that do not rely heavily on large memory blocks. Many of the kernels in our dataset are computationally intensive and do not require significant amounts of block RAM, which makes LUT, DSP, and FF more relevant for the analysis.
We only compare our ForgeHLS with the previous DB4HLS dataset because both ForgeHLS and DB4HLS contain code with directly inserted pragmas, making them suitable for LLM training. In contrast, other datasets typically define pragmas in separate scripts or files, which results in a multi-file structure that complicates usage and is less suitable for efficient LLM training. Additionally, none of the previous datasets provide the ADB graph files required by [2], so we choose to compare our dataset with both the dataset from [2] and the PolyBench portion of ForgeHLS for comparison. Table III exhibits the performance of QoR predictions on GNNs and LLMs.
GNN Performance in the Standard Test. For the Standard Test, the MAPE on #[2] is lower than on ForgeHLS, which can be attributed to two main factors. First, the training and test sets in #[2] lack pragma insertion and consist of only 18,600 kernels, compared to the 429,276 designs in ForgeHLS with diverse pragma variations. The limited diversity in #[2] enables the GNN trained on this dataset to perform better on the standard test set, while the GNN trained on ForgeHLS struggles due to the larger and more complex dataset. Second, to ensure a fair comparison, we trained the GNN model on ForgeHLS with the same configuration, including graph layers, graph dimensions and training epochs. This configuration may have allowed the GNN to capture most of the knowledge from the simpler #[2] dataset, but for the more complex ForgeHLS, larger graph dimensions and more training epochs may improve performance on the standard test set. These results highlight the increased complexity and richness of the code patterns in our dataset, as well as the limitations of GNNs when applied to large, complex datasets. The answer to the first question is that GNNs, while effective in certain scenarios, lack the precision to consistently predict true QoR values for a dataset as extensive as ours. Thus, this leads to the second LLMs direct prediction question.
Generalization Ability: GNN vs. LLM. We evaluate the generalization ability of different trained models on the fair Unseen Test, where LLMs generally show superior performance. This can be attributed to two key factors. First, in terms of the number of learnable parameters, the GNN model in our configuration has approximately 450,000 parameters, while the LoRA fine-tuned LLMs involve over 1,000,000 learnable parameters. This significant difference in capacity indicates that LLMs have a greater ability to capture complex patterns, which is beneficial for accurate QoR prediction on the challenging Unseen Test. Second, LLMs are pretrained on vast amounts of diverse data, which enhances their generalization ability. This extensive pretraining allows LLMs to better handle unseen kernel structure, including analyzing diverse HLS code encountered in the Unseen Test, or capturing pragma information and influence effectively.
Standard vs. Unseen Test. From the results across different datasets, we observe that models generally perform better on the Standard Test Set, as expected, due to its in-distribution nature, meanwhile the Unseen Test Set is more challenging. The Standard Test Set represents the model’s in-distribution fitting ability, where the training and test sets may contain overlapping kernels. In contrast, the Unseen Test Set consists of kernels that are entirely excluded from the training process, providing a stricter evaluation of generalization. The performance drop on the Unseen Test Set emphasizes the challenges of generalizing to new, unseen kernels, especially for GNN models. In contrast, LLMs consistently demonstrate better performance on the Unseen Test Set, indicating their superior ability to generalize across diverse HLS code patterns. This advantage is likely due to their large-scale pretraining and their capacity to capture complex relationships in data.
The QoR prediction task underscores the potential of LLMs in accurately predicting HLS performance, while also presenting new challenges for training GNN prediction models on large-scale datasets.
IV-C Downstream Task 2: Automatic HLS Pragmas Exploration

### v3

@@ 3.1 Dataset Overview
The data construction flow involves two key stages: collection and generation of kernels (Sec. 3.2) and design space exploration for pragma insertion (Sec. 3.3). Based on this flow, we can construct a dataset that is both comprehensive and versatile, suitable for multiple HLS-related downstream tasks (Sec. 4). As shown in Table 2, ForgeHLS significantly outperforms all existing datasets in both scale and complexity. It includes 846 kernels and 459,850 designs, far exceeding prior datasets in volume. More importantly, ForgeHLS offers the highest overall complexity among all datasets, with an average of 95.3 pragmas and 718.6 tokens per kernel. This is driven by its diverse and realistic design sources, especially in R-ForgeHLS, where Vitis-Library reaches up to 244.0 pragmas and CHStone reaches up to 8985.8 tokens per kernel.
Moreover, as illustrated in Figure 2, ForgeHLS demonstrates a broad and dense distribution across multiple post-HLS resource metrics, including LUT, DSP, FF, and BRAM_18K utilization, as well as performance and structural attributes such as worst-case latency, best-case latency, pragma number, and code length. Unlike prior datasets that suffer from narrow, skewed, or fragmented distributions, ForgeHLS provides abundant and continuous coverage across both low and high regimes of these dimensions. This rich and diverse design space facilitates the training of more robust and generalizable models for various HLS learning tasks.
In terms of storage size, ForgeHLS occupies 2177 GB, which is significantly larger than previous datasets. This is primarily due to the inclusion of extensive post-HLS artifacts, particularly graph-based representations such as Control Data Flow Graphs (CDFGs). These graphs are essential for enabling GNN-based learning, making ForgeHLS especially valuable for tasks like QoR prediction, HLS optimization, and performance modeling in ML research.
3.2 Collection and Generation of Kernels
| MAPE | RMSE | ||||||||||||
| Standard Test | Unseen Test | Standard Test | Unseen Test | ||||||||||
| Base Model | Train set | DSP | FF | LUT | DSP | FF | LUT | DSP | FF | LUT | DSP | FF | LUT |
| RGCN | GNN Baseline | 0.1935 | 0.2957 | 0.3197 | >10 | >10 | >10 | 0.0008 | 0.0005 | 0.0008 | >10 | >10 | >10 |
| S-ForgeHLS | 0.5363 | 0.7683 | 0.4071 | 1.1822 | 1.5382 | 0.5736 | 0.0082 | 0.0020 | 0.0029 | 0.0710 | 0.0323 | 0.8404 | |
| ForgeHLS | 0.6038 | 0.6188 | 0.4412 | 0.8728 | 2.7966 | 0.6250 | 0.0181 | 0.0341 | 0.0102 | 0.1369 | 0.0303 | 0.0395 | |
| SAGE | GNN Baseline | 0.1146 | 0.0773 | 0.1678 | 1.3615 | 8.1824 | 1.0354 | 0.0003 | 0.0002 | 0.0002 | 0.0901 | 0.0489 | 0.1012 |
| S-ForgeHLS | 0.5903 | 0.6719 | 0.2800 | 4.0069 | 1.1220 | 0.5368 | 0.0082 | 0.0021 | 0.0026 | 0.6072 | 0.0788 | 0.5015 | |
| ForgeHLS | 0.5448 | 0.7654 | 0.3881 | 0.9652 | 2.3089 | 0.5441 | 0.0096 | 0.0106 | 0.0090 | 0.0790 | 0.0204 | 0.0417 | |
We adopt a dual approach to kernel collection: aggregating real-world HLS kernels from open-source repositories and generating diverse algorithmic implementations using GPT-4o. Representative algorithms are shown in Figure 3.
For the real-world component, we aggregate kernels from previous well-established benchmarks, including:
PolyBench: A set of polyhedral benchmark tests designed to evaluate computational kernels.
MachSuite: A comprehensive suite of machine learning algorithms for hardware synthesis evaluation.
CHStone: A collection of cryptographic algorithms widely used for HLS research and FPGA design.
Rosetta: A benchmark suite containing computer vision algorithms for performance test in hardware accelerators.
Vitis Examples: A repository of FPGA development examples provided by Xilinx’s Vitis HLS toolchain.
Vitis Library: A collection of performance-optimized libraries designed to accelerate various applications on AMD platforms.
Despite leveraging these diverse well-established sources, the Real-world part of ForgeHLS dataset encompassed only 405 kernels. To extend the dataset’s breadth and enhance its applicability, we searched and collected a diverse set of HLS-compatible algorithm specifications list:
HLS Algorithms: Includes matrix multiplication, FFT, signal processing algorithms and deep learning kernels (e.g., LSTM, DNN layers), optimized for hardware.
Operators: Low-level hardware components like adders, counters, and dividers.
Leetcode Algorithms: Representative C/C++ algorithms such as binary search and merge sort, commonly used for algorithmic training.
RTL Algorithms: A specification list of algorithms based on RTL chip, IP core, and module, as described in the RTL dataset (Li et al. 2025).
Subsequently, We prompt GPT-4o with the algorithm specifications and HLS design rules, which generates the corresponding kernel code in C, devoid of pragmas. Then we run the Vitis HLS flow to validate syntactic correctness of our synthetic HLS code.
3.3 Automated Pragma Insertion for Design Generation
After collecting kernel code, we perform syntactic C code analysis based on Clang (Fandrey 2010) and LLVM (Lattner and Adve 2004), and extract essential information needed for DSE, including loop trip counts, loop locations, array sizes, array definition locations, and the identification and location of the top function, as shown in Figure 1.
Full DSE. We design a tree-based explorer (Algorithm 1) to systematically traverse pragma configurations for each kernel. While the traversal is near-complete, it applies HLS-specific constraints to prune invalid or redundant designs—for example, inner loops are not unrolled if the outer loop is pipelined.
However, the design space grows exponentially with the number of pragmas, making full DSE infeasible for complex kernels. For kernels with over 10 pragmas, it can produce millions of candidates, rendering exhaustive HLS synthesis computationally impractical.
Bayesian DSE. To mitigate the combinatorial explosion in large kernels, we propose a Bayesian design space exploration strategy (Algorithm 2). Instead of enumerating all pragma combinations, Bayesian DSE iteratively selects promising configurations using a surrogate model guided by prior synthesis results. Given a kernel , we first extract its pragma space via Clang-based static analysis. The algorithm initializes exploration with randomly sampled configurations. In each iteration, a new design is proposed, synthesized, and evaluated; the result is then used to update the Bayesian optimizer and guide subsequent pragma selection.
Unlike conventional DSE approaches that target Pareto-optimal designs, our objective is to maximize coverage of diverse, high-quality configurations. All valid designs generated during this process are retained and included in the final ForgeHLS dataset.
We conducted over a month of DSE on a high-performance server equipped with 3 nodes, each having 128 CPU cores, totaling over 200k CPU hours. We are confident that our dataset will provide significant value and contribute meaningfully to the research community.
| MAPE | RMSE | ||||||||||||
| Standard Test | Unseen Test | Standard Test | Unseen Test | ||||||||||
| Base Model | Train set | DSP | FF | LUT | DSP | FF | LUT | DSP | FF | LUT | DSP | FF | LUT |
| Deepseek-v3-671B | Zero-shot | - | - | - | 3.6083 | 6.6303 | 1.4584 | - | - | - | 0.2329 | 0.0253 | 0.0926 |
| LLaMA3-7B | Zero-shot | - | - | - | 1.4292 | >10 | 2.2131 | - | - | - | 0.0931 | 0.0401 | 0.0527 |
| DB4HLS | 0.7238 | 9.0168 | 6.0243 | 0.7014 | >10 | >10 | 0.0318 | 0.0044 | 0.0236 | 0.0931 | 0.0401 | 0.0529 | |
| HLSyn | 8.8117 | 1.4946 | >10 | >10 | >10 | >10 | 4.2514 | 0.3345 | 0.8270 | 1.6102 | 0.0426 | 0.2521 | |
| HLSDataset | 1.2517 | 1.2488 | 0.7571 | 1.0063 | >10 | 2.6166 | 0.0138 | 0.0106 | 0.0158 | 0.0918 | 0.0398 | 0.0523 | |
| HLSFactory | 1.2451 | 1.7471 | 1.5141 | >10 | >10 | >10 | 0.0583 | 0.0144 | 0.0417 | 2.6372 | 0.9342 | 1.8684 | |
| S-ForgeHLS | 0.3135 | 0.3565 | 0.2722 | 4.0751 | 1.5098 | 0.7791 | 0.3361 | 0.0305 | 0.1312 | 0.0921 | 0.0399 | 0.0516 | |
| ForgeHLS | 0.2307 | 0.6174 | 0.2604 | 0.5511 | 1.1123 | 0.5970 | 0.0821 | 0.0365 | 0.0689 | 0.0936 | 0.0410 | 0.0525 | |
| Qwen2.5-7B | Zero-shot | - | - | - | >10 | >10 | 4.1366 | - | - | - | 1.7906 | 0.0385 | 0.0536 |
| DB4HLS | 0.8117 | 4.8090 | 3.9484 | 0.7322 | >10 | 9.5164 | 0.0576 | 0.0061 | 0.0253 | 0.0692 | 0.0379 | 0.0653 | |
| HLSyn | 9.1322 | 3.3357 | >10 | 1.1127 | 7.5602 | 2.3977 | 2.0225 | 0.1488 | 0.4875 | 0.1351 | 0.0399 | 0.0532 | |
| HLSDataset | 1.3906 | 1.3399 | 1.0738 | 1.3438 | >10 | 2.1044 | 0.0139 | 0.0109 | 0.0156 | 0.0917 | 0.0397 | 0.0521 | |
| HLSFactory | 0.4442 | 0.5438 | 0.5316 | 9.2722 | >10 | 3.2171 | 0.0946 | 0.0145 | 0.0787 | 0.0946 | 0.0407 | 0.0533 | |
| S-ForgeHLS | 0.1230 | 0.1803 | 0.1503 | 1.0593 | 1.1881 | 0.4918 | 0.3181 | 0.0294 | 0.1189 | 0.0920 | 0.0399 | 0.0521 | |
| ForgeHLS | 0.4666 | 0.4937 | 0.2703 | 1.0150 | 0.8406 | 0.4960 | 0.3394 | 0.0434 | 0.1444 | 0.1528 | 0.0399 | 0.0530 | |
| Mistral-7B-v0.2 | Zero-shot | - | - | - | 8.7915 | 1.0384 | 0.9897 | - | - | - | 0.2032 | 0.0399 | 0.0524 |
| DB4HLS | >10 | >10 | >10 | >10 | >10 | >10 | 2.8257 | 0.0321 | 0.2167 | 4.9136 | 0.0534 | 0.1010 | |
| HLSyn | 3.1403 | 1.2435 | 1.2034 | 7.6851 | >10 | >10 | 2.6475 | 0.2191 | 0.5880 | 0.2581 | 0.0548 | 0.1127 | |
| HLSDataset | 3.1403 | 1.2435 | 1.2034 | 7.6851 | >10 | >10 | 2.6475 | 0.2191 | 0.5880 | 0.2581 | 0.0548 | 0.1127 | |
| HLSFactory | 1.0000 | 0.9758 | 2.1080 | 0.9986 | >10 | >10 | 0.1389 | 0.0256 | >10 | 0.0906 | 7.5828 | >10 | |
| S-ForgeHLS | 0.1141 | 0.1682 | 0.1113 | 0.9039 | 0.6486 | 0.4580 | 0.1219 | 0.0132 | 0.0492 | 0.0919 | 0.0398 | 0.0518 | |
| ForgeHLS | 0.5330 | 0.4575 | 0.2088 | 0.7110 | 0.9706 | 0.5881 | 0.1183 | 0.0258 | 0.0343 | 0.0917 | 0.0398 | 0.0519 | |
4 Evaluation
4.1 Downstream Task 1: HLS QoR Prediction
@@
In this downstream task, we train GNN models and LLMs on various datasets for post-HLS QoR prediction. Specifically, we aim to explore two key questions: 1) Can GNNs capturing the complex patterns within a large-scale dataset like ours? 2) Are LLMs viable for regression tasks in QoR prediction after fine-tuning?
Standard Test and Unseen Test. To evaluate model generalization across different train set fairly, we construct an Unseen Test Set by selecting 10% of kernels from ForgeHLS at the kernel level, resulting in 57 kernels along with all their corresponding designs. This ensures that no kernel in the training set overlaps with those in the Unseen Test Set, enabling a fair assessment of generalization across dataset. In parallel, similar to existing QoR prediction works, we define a Standard Test Set for each dataset by performing a 9:1 split of designs at the design level. Unlike the Unseen Test Set, the Standard Test Set may include designs derived from the same kernels used in training, but with different pragma configurations. It is important to note that the Unseen Test Set does not have any bias towards ForgeHLS, because the kernels using in training set in ForgeHLS is the remaining kernel portion not included in the Unseen Test Set.
Experimental Setup. We adopt the knowledge-rich approach from (Wu et al. 2022) for GNN training, where ADB files are compiled into CDFGs for input. The GNN model consists of 5 layers with a hidden dimension of 300. For LLMs, we fine-tune mainstream open-source models using LoRA (Hu et al. 2021) with Code-QoR pairs as input, as shown in figure 1. The LoRA rank is set to 32, and the LoRA alpha is set to 64. All experiments are conducted on 8 A800 GPUs, using LlamaFactory (Zheng et al. 2024) for fine-tuning and vLLM (Kwon et al. 2023) for inference.
GNN Performance. Table 3 shows the QoR prediction performance of GNN models. On the Standard Test, the GNN Baseline outperforms ForgeHLS in terms of MAPE and RMSE, but it struggles on the Unseen Test, where ForgeHLS performs better. The GNN Baseline, trained on a smaller dataset with only 18,600 kernels, overfits to the simpler patterns of the Standard Test and fails to generalize to new kernels. On the other hand, ForgeHLS performs better on the Unseen Test, suggesting that the graph-based data in ForgeHLS aids generalization. However, on the Standard Test, ForgeHLS underperforms compared to the GNN Baseline, indicating that the complexity of ForgeHLS makes it harder for the GNN to fit the seen kernels. The answer to the first question is that GNNs, while effective in certain scenarios, lack the precision to consistently predict true QoR values for a dataset as extensive as ours. Thus, this leads to the second LLMs direct prediction question.
LLM Performance. Table 4 shows the QoR prediction performance of various LLMs. In zero-shot evaluations, Deepseek-v3-671B performs the best among the LLMs tested. However, after fine-tuning, open-source models significantly outperform Deepseek-v3-671B. For instance, LLaMA3-7B fine-tuned on ForgeHLS achieves a DSP MAPE of 0.5511, a notable improvement over Deepseek’s zero-shot performance of 3.6803. In the Standard Test, S-ForgeHLS exhibits the lowest MAPE (best performance), followed by ForgeHLS. On the Unseen Test, ForgeHLS consistently outperforms other datasets in terms of MAPE, while S-ForgeHLS shows slightly lower performance. Moreover, LLM predictions generally drop on the Unseen Test. For example, Qwen2.5-7B trained on ForgeHLS predicts a LUT MAPE of 0.2703 on the Standard Test, which increases to 0.4960 on the Unseen Test, though still lower than the MAPE of 0.5368 observed for GNN on the Unseen Test. Therefore, the answer to question 2 is that LLMs, after fine-tuning, are indeed viable for regression tasks in QoR prediction, offering significant improvements and better generalization over zero-shot models.
4.2 Downstream Task 2: Automatic HLS Pragmas Exploration